OpenStudy (bekkah323):

A long-distance trucker traveled 120 mi in one direction during a snowstorm. The return trip in rainy weather was accomplished at double the sped and took 3 hr less time. Find the speed going. Using D=rt

OpenStudy (anonymous):

Let X be the speed used to travel there. Time to travel there = Distance / Speed = 120 / X Thus return trip speed = 2X. Time to travel back = Distance / Speed = 120 / 2X Time to travel back = Time to travel there - 3 120 / 2X = 120 / X - 3 120 / 2X = 360 / 3X - 3X / 3X 120 / 2X = 360 - 3X / 3X 120 (3X) = (360 - 3X) (2X) 360X = 720X - 6X^2 6X^2 - 36X = 0 X^2 - 6X = 0 (X)(X-6) = 0 X = ? AND ?

OpenStudy (bekkah323):

is their anyway you could show me how to do it using the D=RT formula?

OpenStudy (anonymous):

D=RT, R is speed, T is time For travelling there: 120 = RT For travelling back: 120 = (2R)(T-3)

OpenStudy (anonymous):

D=RT, R is speed, T is time For travelling there: 120 = RT ---(1) For travelling back: 120 = (2R)(T-3) -----(20 Equate 1 and 2: RT = (2R)(T-3) RT = 2RT - 6R RT - 6R = 0 R(T-6)=0 R=0 or T-6=0 R=0 (N.A), thus T =6

OpenStudy (bekkah323):

so to find r i would just plug in T into 120=(2r)(T-3)

OpenStudy (bekkah323):

@dauspex

OpenStudy (anonymous):

Yes

OpenStudy (bekkah323):

and what would be the first original equation before you start adding the properties of the question? Incase i have to do another problem like this

OpenStudy (anonymous):

D=RT This is the general equation. What i'm doing is to plug in values of D,R,T, if available, then equate them together to try to remove the variables / simplify

OpenStudy (bekkah323):

ahhh ok, thank you